# Orbitronics in Two-dimensional Materials

### v1

@@ IV Orbital Hall effect in 2D materials
The quantum spin Hall insulator is a topological phase where the bulk has an energy band gap indexed by a $\mathbb{Z}_{2}$ topological invariant, which remains robust against smooth changes to the Hamiltonian. When the Fermi energy lies within this gap, linear response theory predicts a finite spin Hall conductivity, a hallmark of its non-trivial topology. Furthermore, creating boundaries in such systems reveals spin-polarized metallic edge states that carry the spin-Hall current, protected by time-reversal symmetry. This concept, introduced by Kane and Mele, is foundational in the study of topological insulators [9].
@@
While these properties are well-established in the spin sector, analogous phenomena in the orbital sector have distinct differences. Refs. [45, 46] demonstrated that a finite OHC plateau can arise in certain insulating phases of 2D systems (see Fig. 3b for the OHC of a MoS2 bilayer). The $p_{x}$-$p_{y}$ model on a honeycomb lattice, used to describe group-V materials grown on SiC substrates [129, 130], serves as an example [46]. In this model, the $p_{z}$ orbital is excluded from the low-energy sector, leaving $p_{x}$ and $p_{y}$ orbitals to dominate electronic transport. Such orbital filtering provides an example of manipulating the OAM nature of a solid by altering the crystalline environment, a possibility widely explored in 2D materials.
@@
Monolayers of TMDs with a 2H structural phase exhibit broken inversion symmetry, as can be seen in Fig. 3a if one considers only one layer. This leads to unique orbital-polarized edge states that cross the band-gap region when shaped into a nanoribbon with zigzag edges [131, 54], shown in Fig. 3c. These edge states are responsible for transporting the OHC [54]. Notably, monolayers of 2H-TMDs also display an OHC plateau within the insulating gap [45]. To further understand the origin of the OHC plateau and to differentiate it from effects purely related to inversion symmetry breaking, the OHC was calculated in bilayers of TMDs, which maintain inversion symmetry (see 3a) [48] and the plateau OHE persists. However, unlike the quantum spin Hall insulator, the orbital Hall insulator cannot be indexed by a $\mathbb{Z}_{2}$ topological invariant. Instead, the orbital Chern number, introduced in Refs. [48, 49], serves as a topological invariant for the orbital Hall insulating phase of 2H-TMDs. Monolayers of 2H-TMDs are characterized by an orbital Chern number $C^{\rm 1l}_{L_{z}}=1$, while bilayers have $C^{\rm 2l}_{L_{z}}=2$. In the low-energy regime, where the Hamiltonian can be written as Dirac-like Hamiltonian, the OHC plateau within the intra-atomic approximation is proportional to the orbital Chern number, $\sigma^{L_{z}}_{xy}=\left(\frac{e}{2\pi}\right)2C_{L_{z}}$ [49]. The edge states in these systems, while related to the orbital Chern number, do not enjoy the same topological protection as those in the quantum spin Hall effect due to the absence of a $\mathbb{Z}_{2}$ invariant. Subsequent studies have further explored this topological invariant in other TMDs, showing their susceptibility to topological phase transitions [120, 121].
@@
In Refs. [50, 51], it was proposed that monolayers of 2H-TMDs exhibit a higher-order topological insulator (HOTI) phase, characterized by the presence of corner states with fractional charge when the layer is fabricated in a flake geometry that preserves threefold rotational symmetry. These corner states manifest as discrete energy levels within the bulk band gap of the energy spectrum, indicating their topologically protected nature and distinguishing them from bulk or edge states (see Figure 3d). This higher-order topology is closely connected to the multi-orbital nature of 2H-TMDs. Subsequently, a study [47] linked this HOTI phase to a finite orbital Hall effect within the insulating gap in numerous TMD compounds that crystallize in two distinct structural phases: the non-centrosymmetric 2H phase and the centrosymmetric 1T phase. The connection between HOTI and OHE in these TMD phases was justified using symmetry arguments and supported by a study with an effective Bernevig-Hughes-Zhang (BHZ) toy model. Subsequent works have extended the connections between topology and the multi-orbital nature of solids, particularly between HOTI and OHE [111, 132], in materials like 2D ferroelectrics and superconductors [133], as well as in 2D ferromagnets [127]. Additionally, well-established topological phenomena, such as the quantum Hall effect, are being reinterpreted with advances in understanding the orbital angular momentum (OAM) of electrons in solids [134].
@@
Figure 3: (a) Side view of a bilayer of 2H-TMD. $\mathcal{I}$ represents the inversion-symmetry point. (b) The OHC plateau predicted by the low-energy theory for a bilayer of 2H-TMD. The blue curve represents the result obtained using the intra-atomic approximation, while the red curve represents the results derived from the Bloch OMM formulation. The dashed line marks the quantized value of the OHE plateau. (c) The energy spectrum of a MoS2 nanoribbon with zigzag edges reveals OAM-polarized edge states that cross the bulk band and can transport OHC. (d) Energy spectra of triangular nanoflakes of MoS2 with in-gap corner states and the corresponding image of the distribution probability associated with the corner-states wave function. Figures adapted from Refs. [48, 49, 47].
@@ V Controlling Orbital Angular Momentum in 2D Materials
Magnetic 2D materials, such as CrI3, exhibit broken time-reversal symmetry, and an equilibrium orbital magnetization may appear. Bilayers of these materials have the interesting properties of breaking time-reversal ($\mathcal{T}\rightarrow\bm{\times}$) and spatial-inversion ($\mathcal{P}\rightarrow\bm{\times}$) symmetries individually while preserving their combined operation ($\mathcal{P}.\mathcal{T}\rightarrow\bm{\checkmark}$). In this case, it is possible to modify the equilibrium orbital magnetization using an electric field without the flow of charge current; that is, a field-induced modification of orbital magnetization can be achieved [147]. This implies that even in an insulating state, the OAM can be controlled. This effect resembles the OREE but is governed by distinct symmetry properties [60] and a different microscopic mechanism. Along with the OREE, it belongs to the more general class of orbital magnetoelectric effects (see details on nomenclature in Ref. [61]). We note that current-induced orbital magnetization can arise even in higher-order terms in the electric field expansion. This phenomenon corresponds to the non-linear OREE [144, 148]. Experiments reporting this effect in 2D materials can already be found [145]. Such non-linear phenomena are often studied by the light-matter interactiomn community where orbitronics in 2D materials is also under investigation [122, 128, 124].
@@ VI Effects of Disorder
Beyond idealized models, real-world applications of orbitronics require an understanding of the disorder, which is ubiquitous in the large-scale fabrication of any device. Building on the parallelism between spintronics and orbitronics, critical metrics for addressing the performance of orbitronic devices, such as the orbital diffusion lengths and the charge-to-orbital conversion, can be altered and even enhanced by the disorder. In the following, we separate the orbital relaxation and the extrinsic contributions to the orbital Hall effect and recollect recent experimental and theoretical work on these topics.
@@
Following the direct experimental detection of the OHE through MOKE measurements, the experimental data raised additional inquiries about the orbital relaxation and the role of the disorder. Choi et al., using MOKE and orbital torque measurements in Ti and Ti/FM systems, obtained orbital diffusion lengths of $74\pm 24$ nm and $50\pm 15$ nm [31], respectively, with the latter being in agreement with previous estimations on similar heterostructures [73]. In contrast, similar Moke measurements from Lyalin et al. in Cr reported orbital relaxation lengths of $6.6\pm 0.6$ nm [32]. On the same order of magnitude, orbital Hanle measurements from Sala et al. in polycrystalline Mn suggested orbital relaxation lengths of $\sim 2$ nm [33], which are in close agreement with the estimates from Liu et al. in CoPt/Cu/MgO heterostructures [153].
@@
Though these experimental data seem to point out that disorder might be a limiting factor for orbital conductivity, there is experimental data that shows the contrary. Kim et al., using orbital torque measurements, found orbital relaxation lengths in Cu samples that are about 10 nm, in agreement later the estimations from Kim et al. for CoFe/Cu/TiO2 and CoFe/Cu/SiO2 where the O from the oxide hybridizes with the Cu and consequently favors the orbital transport [154, 155]. Along the same lines, experiments from Seifert et al. suggested that ballistic orbital propagation occurs over $20$ nm - in agreement with previous experiments from Hayashi et al. [73] - and little influence of the magnetic impurities on the orbital propagation [79]. Furthermore, recent experiments from Idrobo et al. analyzed the orbital relaxation in Ti-based devices and found orbital relaxation lengths of $7.9$ nm with a strong dependence on sample morphology [156].
The differing estimates of orbital diffusion lengths suggest that orbital dynamics function differently from spin dynamics, offering a potential method to distinguish between spin and orbital signals. On the theoretical front, Han et al. investigated orbital dynamics in centrosymmetric systems. Combining symmetry arguments and quantum kinetic theory, they demonstrated the orbital and spin degrees of freedom have distinctive anticommutation relations that allow for orbital-momentum locking even in centrosymmetric systems. These couplings give rise to the orbital textures (different from OAM textures, as discussed above), contribute to the OAM dynamics, and, in nonequilibrium, give rise to the orbital Hall effect in metals [157].
@@
Sohn et al. investigated the coupled spin-orbital dynamics in centrosymmetric materials using quantum kinetic theory and device simulations [158]. Their study demonstrated that orbital-momentum locking, arising from the orbital texture, leads to a Dyakonov-Perel-like orbital relaxation, suggesting that orbital angular momentum undergoes rapid decoherence. Conversely, Ref. 159 found a different behavior when studying the decay of orbital currents using first-principles quantum mechanical scattering calculations. By injecting orbital currents from an orbitally polarized lead into bulk transition metals, they observed that the orbital current persisted for only a few atomic layers, indicating a different relaxation mechanism than the one predicted in Ref. [158].
@@
Recent calculations for a 2D tight-binding model using a quantum Boltzmann equation approach that takes into account different impurity models show that the impurity symmetry strongly influences the OAM relaxation that can change from Dyakonov-Perel-like to Elliot-Yafet relaxation if the axial symmetry of the impurity is broken [160].
@@
Discrepancies in the estimates of orbital diffusion lengths also open the possibility for the influence of extrinsic contributions that may alter the orbital currents and OAM accumulation. Since the early days of spintronics, the effects of the disorder have been thoroughly studied and discussed [2]. Beyond influencing spin relaxation and decoherence, disorder can also give rise to the SHE through extrinsic mechanisms such as skew scattering and side-jump. These impurity-driven effects can generate spin currents even in materials where intrinsic mechanisms are weak or absent, demonstrating that disorder can play a fundamental role in spin transport phenomena.
@@
In orbitronics, similar extrinsic mechanisms are expected to influence orbital transport, potentially generating OHE. Just as skew scattering and side-jump effects can drive SHE, these mechanisms are predicted to play a role in orbital dynamics by deflecting OAM currents in the presence of impurities. While the study of disorder-driven orbital transport is still in its early stages, recent works have begun to explore how extrinsic contributions affect OHE and orbital accumulation. Furthermore, the dichotomy between the atom-centered (intra-site) approximation and itinerant contributions (inter-site) for the orbital moment raises even more questions about the effects of the disorder on the orbital moment generation.
@@
In the atomic centered approximation (ACA), the seminal paper from Bernevig et al. discusses the effect of disorder in p-doped semiconductors, where they argued that due to the symmetry of the system, the disorder effects enter as self-energy corrections in their linear response calculations due to the cancellation of the vertex corrections in the current operators [17]. Tanaka et al. demonstrated that for transition metals, the vertex correction considered within the first Born approximation has a negligible impact on the spin and orbital Hall conductivities [19]. Similar trends were obtained in recent work by Pezo et al. [116]. Recent work based on quantum Boltzmann equations suggests the existence of lattice symmetries where disorder might lead to vanishing orbital Hall currents [161]. Conversely, calculations for a 2D tight-binding model, a quantum Boltzmann equation approach with third order of perturbation theory shows the appearance of a sizable extrinsic contribution to the orbital Hall effect that depends on the impurity concentration $n_{i}$ [160].
When it comes to the disorder effects on the OHE of itinerant orbital momentum, there is a set of recent theoretical results in 2D materials. Recent work from Liu et al. using quantum kinetic equations with the perturbative treatment of the disorder pointed out the appearance of extrinsic contributions that rapidly overcome the intrinsic contributions [153]. Subsequently, Canonico et al. developed a real-space numerical methodology capable of capturing disorder effects on the OHC nonperturbatively [102]. Their calculations demonstrated that the trends obtained from the perturbation theory by Liu et al. and beyond the reach of perturbative treatments hinted that disorder can enhance the OHC from the ballistic up to the diffusive regime until the onset of localization where the orbital transport is reduced.
@@
Further theoretical studies by Veneri et al., using a nonperturbative descriptions of disorder in the dilute regime, found that the extrinsic contributions to the OHC are dominated by the skew-scattering mechanism. In this regime, the OHC is proportional to the inverse of the impurity concentration, meaning that as impurity concentration decreases, the skew-scattering effect becomes more significant, leading to an increase in OHC. These findings elucidate the origin of extrinsic contributions and emphasize the role of perturbation symmetries in shaping the extrinsic OHC [115]. A subsequent work performed a similar analysis but for centrosymmetric bilayer 2H-TMD [119].
Regarding the orbital Rashba Edelstein effect, the ACA formulation is almost unexplored. In contrast, the effects of the disorder on the itinerant orbital moments have been studied more extensively. In equilibrium calculations, within Keldysh Green’s function formalism, Zhu et al. obtained a closed form for the equilibrium orbital moment for magnetic disordered systems [162]. While analyzing the magnetic 2DEG Rashba model, they found that white noise disorder does not affect the orbital moment, representing a shift in the redistribution of the orbital moment on energy due to the renormalization of the energy states caused by the disorder. Beyond this, Rou et al., using semiclassical calculations, found that the charge-to-orbital conversion acquires two extrinsic contributions due to the weak disorder: skew-scattering and side-jump [163]. Although weak disorder can enhance charge-to-orbital conversion, due to its linear dependence on the scattering time, it can be rapidly disturbed in heavy disorder situations. Still, further investigation on the non-perturbative disordered regime in disordered 2D materials is required.

### v2

@@ IV Orbital Hall effect in 2D materials
The quantum spin Hall insulator is a topological phase where the bulk of the system has an energy band gap indexed by a $\mathbb{Z}_{2}$ topological invariant, which remains robust against smooth changes to the Hamiltonian. When the Fermi energy lies within this gap, linear response theory predicts a finite spin Hall conductivity, a hallmark of its non-trivial topology. Furthermore, creating boundaries in such systems reveals spin-polarized metallic edge states that carry the spin-Hall current, protected by time-reversal symmetry. This concept, introduced by Kane and Mele, is foundational in the study of topological insulators [9].
@@
While these properties are well-established in the spin sector, their analogs in the orbital sector exhibit pronounced differences. Refs. [45, 46] demonstrated that a finite OHC plateau can arise in certain insulating phases of 2D systems (see Fig. 4b for the OHC of a MoS2 bilayer). The $p_{x}$-$p_{y}$ model on a honeycomb lattice, used to describe group-V materials grown on SiC substrates [132, 133], serves as an example [46]. In this model, the $p_{z}$ orbital is excluded from the low-energy sector, leaving $p_{x}$ and $p_{y}$ orbitals to dominate electronic transport. Such orbital filtering provides an example of manipulating the OAM nature of a solid by altering the crystalline environment, a possibility widely explored in 2D materials.
@@
Monolayers of TMDs with a 2H structural phase exhibit broken inversion symmetry, as can be seen in Fig. 4a if one considers only one layer. This leads to unique OAM-polarized edge states that cross the band-gap region when shaped into a nanoribbon with zigzag edges [134, 54], shown in Fig. 4c. These edge states are responsible for transporting the OHC [54]. Notably, monolayers of 2H-TMDs also display an OHC plateau within the insulating gap [45]. To further understand the origin of the OHC plateau and to differentiate it from effects purely related to inversion symmetry breaking, the OHC was calculated in bilayers of TMDs, which maintain inversion symmetry (see 4a) [48] and the plateau OHE persists. However, unlike the quantum spin Hall insulator, the orbital Hall insulator cannot be indexed by a $\mathbb{Z}_{2}$ topological invariant. Instead, the orbital Chern number, introduced in Refs. [48, 49], serves as a topological invariant for the orbital Hall insulating phase of 2H-TMDs. Monolayers of 2H-TMDs are characterized by an orbital Chern number $C^{\rm 1l}_{L_{z}}=1$, while bilayers have $C^{\rm 2l}_{L_{z}}=2$. In the low-energy regime, where the Hamiltonian of 2H-TMDs can be written as Dirac-like Hamiltonian, the OHC plateau within the intra-atomic approximation is proportional to the orbital Chern number, $\sigma^{L_{z}}_{xy}=\left(\frac{e}{2\pi}\right)2C_{L_{z}}$ [49]. The edge states in these systems, while related to the orbital Chern number, do not enjoy the same topological protection as those in the quantum spin Hall effect due to the absence of a $\mathbb{Z}_{2}$ invariant. Subsequent studies have further explored this topological invariant in other TMDs, showing their susceptibility to topological phase transitions [123, 124].
@@
In Refs. [50, 51], it was proposed that monolayers of 2H-TMDs exhibit a higher-order topological insulator (HOTI) phase, characterized by the presence of corner states with fractional charge when the layer is fabricated in a flake geometry that preserves threefold rotational symmetry. These corner states manifest as discrete energy levels within the bulk band gap of the energy spectrum, indicating their topologically protected nature and distinguishing them from bulk or edge states (see Figure 4d). This higher-order topology is closely connected to the multi-orbital nature of 2H-TMDs. Subsequently, a study [47] linked this HOTI phase to a finite orbital Hall effect within the insulating gap in numerous TMD compounds that crystallize in two distinct structural phases: the non-centrosymmetric 2H phase and the centrosymmetric 1T phase. The connection between HOTI and OHE in these TMD phases was justified using symmetry arguments and supported by a study with an effective Bernevig-Hughes-Zhang (BHZ) toy model. Table 1 shows the transverse orbital conductivity of TMD insulating monolayers crystallized in the 1T and 2H structural phases. The structure of each material was obtained from the C2DB database [135]. The transverse spin conductivity vanishes within the energy bandgap for all compounds listed in the tables. Subsequent works have extended the connections between topology and the multi-orbital nature of solids, particularly between HOTI and OHE [114, 136], in materials like 2D ferroelectrics and superconductors [137], as well as in 2D ferromagnets [130]. Additionally, well-established topological phenomena, such as the quantum Hall effect, are being reinterpreted with advances in understanding the orbital angular momentum (OAM) of electrons in solids [138].
2H-TMD $\sigma^{L_{y}}_{yx}$ $\sigma^{L_{z}}_{yx}$ $||\sigma^{L}_{yx}|$$|\sigma^{S}_{yx}|$ $E_{g}(\text{eV})$
MoS2 0.000.000.00 2.652.652.65 2.652.652.65 0.000.000.00 1.601.601.60
MoSe2 0.000.000.00 2.632.632.63 2.632.632.63 0.000.000.00 1.341.341.34
MoTe2 0.000.000.00 2.472.472.47 2.472.472.47 0.000.000.00 0.950.950.95
CrS2 0.000.000.00 1.891.891.89 1.891.891.89 0.000.000.00 0.900.900.90
CrSe2 0.000.000.00 1.921.921.92 1.921.921.92 0.000.000.00 0.710.710.71
CrTe2 0.000.000.00 1.911.911.91 1.911.911.91 0.000.000.00 0.470.470.47
WS2 0.000.000.00 1.431.431.43 1.431.431.43 0.000.000.00 1.561.561.56
WSe2 0.000.000.00 1.631.631.63 1.631.631.63 0.000.000.00 1.271.271.27
WTe2 0.000.000.00 1.461.461.46 1.461.461.46 0.000.000.00 0.770.770.77
TiS2 0.000.000.00 0.780.780.78 0.780.780.78 0.000.000.00 0.720.720.72
TiSe2 0.000.000.00 0.750.750.75 0.750.750.75 0.000.000.00 0.530.530.53
1T-TMD $\sigma^{L_{y}}_{yx}$ $\sigma^{L_{z}}_{yx}$ $||\sigma^{L}_{yx}|$$|\sigma^{S}_{yx}|$ $E_{g}(\text{eV})$
NiS2 1.021.021.02 1.78-1.78- 1.78 2.052.052.05 0.000.000.00 0.540.540.54
PdS2 0.730.730.73 1.22-1.22- 1.22 1.421.421.42 0.00 1.141.141.14
PdSe2 0.850.850.85 1.65-1.65- 1.65 1.851.851.85 0.000.000.00 0.520.520.52
PtTe2 0.820.820.82 1.59-1.59- 1.59 1.701.701.70 0.000.000.00 0.370.370.37
PtSe2 0.600.600.60 1.01-1.01- 1.01 1.181.181.18 0.000.000.00 1.171.171.17
PtS2 0.510.510.51 0.77-0.77- 0.77 0.920.920.92 0.000.000.00 1.721.721.72
ZrS2 0.610.610.61 0.34-0.34- 0.34 0.700.700.70 0.000.000.00 1.161.161.16
ZrSe2 0.840.840.84 0.58-0.58- 0.58 1.021.021.02 0.000.000.00 0.340.340.34
HfS2 0.320.320.32 0.08-0.08- 0.08 0.330.330.33 0.000.000.00 1.271.271.27
HfSe2 0.710.710.71 0.40-0.40- 0.40 1.021.021.02 0.000.000.00 0.480.480.48
[…]
Table 1: The heights of the transversal OAM conductivity plateaus for insulating TMD monolayers, which crystallize in the non-centrosymmetric 2H (shown in the left table) and centrosymmetric 1T (shown in the right table) structural phases. The first column displays the stoichiometric formulas for TMD compounds. In the second and third columns, it shows the values of $\sigma^{L_{y}}_{yx}$ and $\sigma^{L_{z}}_{yx}$ in units of $e/2\pi$. For both structural phases, $\sigma^{L_{x}}_{yx}=0$. The fourth and fifth columns show the values of $|\sigma^{S}_{yx}|$ and $|\sigma^{L}_{yx}|$ in unities of $e/2\pi$, where $|\sigma^{X}_{yx}|=\sqrt{(\sigma^{X_{x}}_{yx})^{2}+(\sigma^{X_{y}}_{yx})^{2}+(\sigma^{X_{z}}_{yx})^{2}}$ for $X=L$ and $S$. The last column presents the energy bandgap for each compound. Details of the numerical calculations of conductivities, along with information on the higher-order topological indexing of each compound, can be found in the supplementary material of Ref. [47]. 1T-TMDs are HOTIs characterized by $\mathbb{Z}_{4}=2$ mod 444 topological indicators. 2H-TMDs are also HOTIs but they are characterized by corner states with a charge $Q^{(3)}_{c}=2e/3$ mod $e$. (∗) NiO2, PdO2, PtO2 are not TMDs, but they share the same structural, electronic, and topological properties.
@@
Figure 4: Aspects of electronic structure, lattice symmetry, OAM transport, and topology of 2H-TMDs. (a) Side view of a bilayer of 2H-TMD. $\mathcal{I}$ represents the inversion-symmetry point. (b) The OHC plateau predicted by the low-energy theory for a bilayer of 2H-TMD. The blue curve represents the result obtained using the intra-atomic approximation, while the red curve represents the results derived from the Bloch OMM formulation. The dashed line marks the quantized value of the OHE plateau. (c) The energy spectrum of a MoS2 nanoribbon with zigzag edges reveals OAM-polarized edge states that cross the bulk band and can transport OHC. (d) Energy spectra of triangular nanoflakes of MoS2 with in-gap corner states and the corresponding image of the distribution probability associated with the corner-states wave function. Figures adapted from Refs. [48, 49, 47].
@@ V Controlling Orbital Angular Momentum in 2D Materials
Magnetic 2D materials, such as CrI3, exhibit broken time-reversal symmetry, and an equilibrium orbital magnetization may appear. Bilayers of these materials with interlayer antiferromagnetic configuration have the interesting properties of breaking time-reversal ($\mathcal{T}\rightarrow\bm{\times}$) and spatial-inversion ($\mathcal{P}\rightarrow\bm{\times}$) symmetries individually while preserving their combined operation ($\mathcal{P}.\mathcal{T}\rightarrow\bm{\checkmark}$). In this case, it is possible to modify the equilibrium orbital magnetization using an electric field without the flow of charge current; that is, a field-induced modification of orbital magnetization can be achieved [151]. This implies that even in an insulating state, the OAM can be controlled. This effect resembles the OREE but is governed by distinct symmetry properties [62] and a different microscopic mechanism. Along with the OREE, it belongs to the more general class of orbital magnetoelectric effects (see details on nomenclature in Ref. [63]). We note that induced orbital magnetization can arise even in higher-order terms in the electric field expansion. This phenomenon corresponds to the non-linear orbital Edelstein effect [148, 152]. Experiments reporting this effect in 2D materials are already available [149]. Such non-linear phenomena are often studied by the light-matter interaction community, where orbitronics in 2D materials is also under investigation [125, 131, 127].
@@ VI Effects of Disorder
Beyond idealized models, real-world applications of orbitronics require further investigation into the role of disorder, which is prevalent in the large-scale fabrication of any device. Building on the parallelism between spintronics and orbitronics, critical metrics for addressing the performance of orbitronic devices, such as orbital diffusion lengths and charge-to-orbital conversion efficiency, can be altered and even enhanced by disorder. In the following, we separate the orbital relaxation and the extrinsic contributions to the orbital Hall effect and recollect recent experimental and theoretical work on these topics.
@@
The experimental values of the orbital diffusion length obtained by different methods and for different materials vary considerably, prompting further investigations into orbital relaxation and the influence of disorder. For example, Choi et al. [31], using MOKE and orbital torque measurements in Ti and Ti/FM systems, obtained orbital diffusion lengths of 74±2474$\pm$ 2474 ± 24 nm and 50±1550$\pm$ 1550 ± 15 nm, respectively. The latter value agrees with previous estimations made for similar heterostructures [75]. In contrast, similar MOKE measurements conducted by Lyalin et al. in Cr reported orbital relaxation lengths of 6.6±0.66.6$\pm$ 0.66.6 ± 0.6 nm [32]. Similarly, orbital Hanle magnetoresistance measurements of Sala et al. in polycrystalline Mn suggested orbital relaxation lengths of approximately 222 nm [33], which, surprisingly, are in close agreement with the estimates of Liu et al. in CoPt/Cu/MgO heterostructures [157].
@@
Although these experimental data suggest that disorder may limit orbital conductivity, other studies indicate the opposite. Kim et al. [158] measured orbital relaxation lengths of approximately 101010 nm in Cu samples using orbital torque measurements. These findings align with later estimations for CoFe/Cu/TiO2 and CoFe/Cu/SiO2 structures, where the oxygen atoms of the oxides hybridize with copper, favoring orbital transport [159]. Similarly, Seifert et al. [82] observed ballistic orbital propagation over distances exceeding 202020 nm, consistent with earlier experiments by Hayashi et al. [75]. Furthermore, recent experiments by Idrobo et al. [160] examined orbital relaxation in Ti-based devices and found orbital relaxation lengths of 7.97.97.9 nm, with a strong dependence on sample morphology.
The varying estimates of orbital diffusion lengths suggest that orbital dynamics operates differently from spin dynamics, which could potentially provide a method to distinguish between spin and orbital signals.
Han et al. [161] noted that spin and orbital angular momentum operators exhibit distinct anti-comutator relations. Specifically, while $\left\{S_{i},S_{j}\right\}\propto\delta_{ij}$, $\left\{L_{i},L_{j}\right\}$ is proportional to the orbital quadrupole density [162]. These characteristics enable ${\bf L}-{\bf k}$ coupling, but prevent ${\bf S}-{\bf k}$ coupling in centrosymmetric non-magnetic systems. As a result, this can lead to unique features in orbital current diffusion that are not present in spin current diffusion, highlighting qualitative differences between orbital and spin dynamics.
@@
Sohn et al. [163] investigated coupled spin-orbital dynamics in centrosymmetric materials using quantum kinetic theory and device simulations. Their study demonstrated that orbital-momentum locking, arising from the orbital texture, leads to a Dyakonov-Perel-like orbital relaxation, suggesting that frequent momentum scattering disrupts orbital precession, reducing dephasing and thereby increasing relaxation times. Conversely, Ref. 164 found a different behavior when studying the decay of orbital currents using first-principles quantum-mechanical scattering calculations. By injecting orbital currents from an orbitally polarized lead into bulk transition metals, they observed that the orbital current persisted for only a few atomic layers, indicating a relaxation mechanism different from the one predicted in Ref. [163].
@@
Recent calculations for a 2D tight-binding model using a quantum Boltzmann equation approach that takes into account different impurity models show that the impurity-potential symmetry strongly influences the OAM relaxation, which can change from Dyakonov-Perel-like to Elliot-Yafet relaxation if the axial symmetry of the impurity potential is broken [165].
Despite growing interest in orbitronics, there are currently no direct measurements or systematic studies of the orbital diffusion length in 2D materials. Most of what is known experimentally concerns spin transport in TMDs, where spin diffusion lengths are typically short. However, orbital relaxation mechanisms can differ substantially. For example, in monolayer 2H-TMDs, the OAM near the ${\bf K}$ and ${\bf K}^{\prime}$ points is strongly constrained by symmetry, which inhibits scattering events that would flip the orbital moment. This symmetry protection suggests the possibility of enhanced orbital diffusion in selected regions of momentum space. While this remains speculative, it motivates further theoretical and experimental work to clarify whether long-range orbital transport can be realized in these systems.
@@
Discrepancies in estimates of orbital diffusion lengths also suggest that extrinsic contributions may influence orbital currents and OAM accumulation.
@@
Since the early days of spintronics, the effects of the disorder have been thoroughly studied and discussed [2]. Beyond influencing spin relaxation and decoherence, disorder can also give rise to the SHE through extrinsic mechanisms such as skew scattering and side-jump. These impurity-driven effects can generate spin currents even in materials where intrinsic mechanisms are weak or absent, demonstrating that disorder can play a fundamental role in spin transport phenomena.
@@
In orbitronics, similar extrinsic mechanisms are expected to influence orbital transport, potentially generating OHE. Just as skew scattering and side-jump effects can drive the SHE, these mechanisms are also predicted to affect orbital dynamics by deflecting OAM currents in the presence of impurities. Although the study of disorder-driven orbital transport is still in its early stages, recent works have begun to explore how extrinsic contributions affect the OHE and orbital accumulation. It is noteworthy that these studies deal with effects of disorder on orbital transport by using either the atom-centered (intra-site) approximation (ACA) or a purely itinerant OAM, which incorporates inter-site contributions.
@@
The seminal paper of Bernevig et al. [17] discusses the effect of disorder in p-doped semiconductors within the ACA. They argued that due to the symmetry of the system under investigation, the effects of disorder appear as self-energy corrections in their linear response calculations, due to cancellation of the vertex corrections in the current operators. Tanaka et al. [19] demonstrated that, for transition metals, the vertex correction considered within the first Born approximation has a negligible impact on the spin and orbital Hall conductivities. Similar trends were obtained in recent work by Pezo et al. [119].
Recent work by Tang and Bauer [166] investigate how random defect scattering influences the OHE. They used a quantum Boltzmann equation to examine a two-band ($p_{x}$-$p_{y}$) model on a triangular lattice, incorporating the so-called vertex corrections. Their findings reveal that even a weak disorder can significantly affect and may completely suppress the intrinsic orbital Hall current.
In contrast, for a 2D tight-binding model, a quantum Boltzmann equation approach with third-order perturbation theory shows the appearance of a sizable extrinsic contribution to the OHE. The extrinsic contribution depends on the impurity concentration, in line with the skew-scattering mechanism [165].
These results show that the influence of vertex corrections on OHE varies depending on the specific model studied. In principle, the discrepancies can be explained with Dimitrova’s argument [167], which was originally introduced in the context of the spin Hall effect. This argument is based on general conservation laws that govern the time evolution of the system [168], which depend on the lattice symmetry and the angular momentum character of the model. For instance, by applying Dimitrova’s argument to orbital transport in the $p_{x}$-$p_{y}$ model on a triangular lattice, one obtains that the orbital Hall conductivity must vanish, which is consistent with the findings of Ref.[166].
Meanwhile other studies explored the OHE in two-dimensional massive Dirac fermions, originating from the itinerant OAM. Liu and Culcer [129] employed a quantum kinetic approach to study the effect of short-range Gaussian disorder, taking into account vertex corrections. They discovered that in doped systems, extrinsic Fermi surface contributions can lead to a substantial increase of the OHE, which also aligns with Dimitrova’s argument. However, their approach does not capture the dependence of the OHE on impurity concentration and strength.
Canonico et al. [105] developed a real-space numerical methodology capable of capturing disorder effects on the OHC non-perturbatively. Their results suggest that disorder can enhance the OHE across different transport regimes—from ballistic to diffusive—and eventually suppress the OHE near localization.
@@
Further theoretical studies by Veneri et al. [118], using a nonperturbative description of disorder in the dilute regime, found that extrinsic contributions to the OHC are dominated by the skew-scattering mechanism. In this regime, the OHC is proportional to the inverse of the impurity concentration, meaning that as impurity concentration decreases, the skew-scattering effect becomes more significant, leading to an increase in OHC. These findings elucidate the origin of extrinsic contributions and emphasize the role of perturbation symmetries in shaping the extrinsic OHC. A subsequent work performed a similar analysis but for centrosymmetric bilayer 2H-TMDs[122].
These findings challenge the prevailing notion that intrinsic mechanisms dominate the OHE, suggesting that extrinsic scattering plays a crucial role, especially for experimentally relevant impurity concentrations.
Most studies examining the effects of disorder on the orbital Rashba-Edelstein effect adopt the Bloch OMM approach. Zhu et al. [169] derived a general formula for calculating the orbital magnetization in disordered systems, using the Keldysh Green’s function formalism. They applied it to calculate the orbital magnetization of a weakly disordered two-dimensional electron gas (2DEG) with Rashba spin-orbit coupling, and found that the orbital magnetization is robust to weak-scattering short-range disorder, which mainly causes a rigid energy shift in the orbital magnetization distribution.
Additionally, Rou et al. [170] found through semiclassical calculations that charge-to-orbital conversion has two extrinsic contributions due to weak disorder: skew-scattering and side-jump. Although weak disorder can enhance charge-to-orbital conversion, due to its linear dependence on the scattering time, it can be rapidly disturbed in heavy disorder situations. Further investigation into the non-perturbative disordered regime in 2D materials is needed.